Tembo Inc 3. The data below is provided for two competing companies Safari Corp Selling Price per unit $40 Variable cost per just do f and g please!
Answer #1

Hello!

Revenue function is Product price per unit multiplied by number of units sold.

i.e. R = P X Q

Cost function is Total Fixed Cost plus Variable Cost multiplied by numner of units.

i.e. C(x) = FC + VC(x)

Break even point is the number of units whose total cost is equal to total revenue, i.e no profit no loss.

As you only want answers for point F and G, so here goes:

Point (f) In the situation of recession, the company sold only 500 units per month. So using the revenue functions and cost functions we will compute the net profit.

For SAFARI CORP,

Total Cost = FC + VC

= 30000 + (10*500)

= 30000 + 5000

= $ 35000

Revenue  = Selling Price X no of units sold

= 40*500

= $ 20000

Net Loss = $ 15000

Similarly for TEMBO INC,

Total Cost = FC + VC

= 10000 + (30*500)

= 10000 + 15000

= $ 25000

Revenue  = Selling Price X no of units sold

= 40*500

= $ 20000

Net Loss = $ 5000

Both the companies suffered losses in recession by selling only 500 units. However, loss of TEMBO INC was lower than loss of SAFARI CORP. Therefore, out of the these two companies, TEMBO INC is better. Also note, TEMBO INC had lower Fixed cost which helped to cop up in recession like scenarios.

Point (g) In the situation of boom, the company sold 3000 units per month. So using the revenue functions and cost functions we will compute the net profit.

For SAFARI CORP,

Total Cost = FC + VC

= 30000 + (10*3000)

= 30000 + 30000

= $ 60000

Revenue  = Selling Price X no of units sold

= 40*3000

= $ 120000

Net profit = $ 60000

Similarly for TEMBO INC,

Total Cost = FC + VC

= 10000 + (30*3000)

= 10000 + 90000

= $ 100000

Revenue  = Selling Price X no of units sold

= 40*3000

= $ 120000

Net Profit = $ 20000

Both the companies had profits in boom by selling 3000 units. However, profits of SAFARI CORP were higher than profits of TEMBO INC. Therefore, out of the these two companies, SAFARI CORP is better. Also note, SAFARI CORP had lower Variable cost (therefore higher margin per unit) which helped to achieve higher profits in boom like scenarios.